# Total variation bounds in the Lindeberg central limit theorem

### v1

@@ 1 Introduction
for all $\varepsilon>0,$ where $\rm 1\hskip-3.87495pt1$} is the indicator function, then $S_{n}$ converges in distribution to a standard normal random variable NN as $n\to\infty.$ Furthermore, if the random variables $X_{k}$ satisfy the Feller-Lévy condition
@@
where $c$ is a positive absolute constant. We now consider the total variation distance between the laws of $S_{n}$ and NN defined by
@@
Corollary $1.$
Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem $1.$ Assume additionally that $\sup\limits_{k\geq 1}J(X_{k})<\infty.$ The, the following two statements are equivalent
@@ 2 Proofs
Lemma $2.$
Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem $1.$ For each $1\leq k\leq n,$ we put
@@
$\rho_{k,n}(x)=\mathbb{E}\left[\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})}{b_{k,n}}\bigg|S_{k,n}=x\right]$ (2.3)
@@
Then, by using the same arguments as in the proof of Proposition 3.2 in [8], we conclude that the law of $S_{k,n}$ has an absolutely continuous density and its score function is given by (2.3). This finishes the proof of the lemma. ∎
@@
Proof of Theorem $1.$ We will carry out the proof in two steps.
@@
Fixed $1\leq k\leq n.$ Let $S_{k,n}$ be as in Lemma $2.$ By the independence of $X_{k}$ and $S_{k,n},$ we can use the relation (2.4) to get
@@
Proof of Corollary $1.$ The implication $(ii)\Rightarrow(i)$ is obvious because of the fact that the convergence in total variation implies the convergence in distribution. To verify the implication $(i)\Rightarrow(ii),$ we observe that, for any $\varepsilon\in(0,1),$ we have

### v2

@@ 1 Introduction
for all $\varepsilon>0,$ where $\rm 1\hskip-3.87495pt1$} is the indicator function, then $S_{n}$ converges in distribution to a standard normal random variable $N$ as $n\to\infty.$ Furthermore, if the random variables $X_{k}$ satisfy the Feller-Lévy condition
@@
where $c$ is a positive absolute constant. We now consider the total variation distance between the laws of $S_{n}$ and $N$ defined by
@@
Corollary 1.1.
Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem 1.1. Assume additionally that $\sup\limits_{k\geq 1}J(X_{k})<\infty.$ The, the following two statements are equivalent
@@
We use the condition $J(X_{k})<\infty$ is to establish the bound (1.6). However, this condition can be removed to recover the classical central limit theorem.
Corollary 1.2.
Let $(Y_{k})_{k\geq 1}$ be a sequence of independent real-valued random variables $\mathbb{E}Y_{k}=0$ and finite variances $\alpha_{k}^{2}=\mathbb{E}Y_{k}^{2}\in(0,\infty).$ Assume that this sequence satisfies Lindeberg’s condition:
$\lim\limits_{n\to\infty}\frac{1}{a_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[Y_{k}^{2}{\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|>\varepsilon a_{n}\}}]=0$ (1.7)
for all $\varepsilon>0,$ where $a_{n}:=\sqrt{\alpha_{1}^{2}+\cdots+\alpha_{n}^{2}}.$ Then, as $n\to\infty,$ we have $\frac{Y_{1}+\cdots+Y_{n}}{a_{n}}\to N$ in distribution.
@@ 2 Proofs
Lemma 2.2.
Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem 1.1. For each $1\leq k\leq n,$ we put
@@
$\rho_{k,n}(x)=\mathbb{E}\left[\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})}{b_{k,n}}\bigg|S_{k,n}=x\right],$ (2.3)
where $\rho_{i}$ denotes the score function of $X_{i}.$
@@
Then, by using the same arguments as in the proof of Proposition 3.2 in our recent paper [8], we conclude that the law of $S_{k,n}$ has an absolutely continuous density and its score function is given by (2.3). This finishes the proof of the lemma. ∎
@@
Proof of Theorem 1.1. We will carry out the proof in two steps.
@@
Fixed $1\leq k\leq n.$ Let $S_{k,n}$ be as in Lemma 2.2. By the independence of $X_{k}$ and $S_{k,n},$ we can use the relation (2.4) to get
@@
Proof of Corollary 1.1. The implication $(ii)\Rightarrow(i)$ is obvious because of the fact that the convergence in total variation implies the convergence in distribution. To verify the implication $(i)\Rightarrow(ii),$ we observe that, for any $\varepsilon\in(0,1),$ we have
@@
Proof of Corollary 1.2. Let $(N_{k})_{k\geq 1}$ be a sequence of independent normal random variables, and independence of $(Y_{k})_{k\geq 1}.$ For each $k\geq 1,$ $\mathbb{E}N_{k}=0$ and $\mathbb{E}N_{k}^{2}=\mathbb{E}Y_{k}^{2}.$ We consider the sequence $X_{k}=Y_{k}+N_{k},k\geq 1$ and define $S_{n}$ as in (1.1). For every $\varepsilon>0,$ we have
$\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]$ $=\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|Y_{k}+N_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|Y_{k}+N_{k}|>\varepsilon b_{n}\}}]$
$\leq\frac{2}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[(|Y_{k}|^{2}+|N_{k}|^{2}){\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|+|N_{k}|>\varepsilon b_{n}\}}]$
$\leq\frac{2}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[(|Y_{k}|^{2}+|N_{k}|^{2})({\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|>\varepsilon b_{n}/2\}}+{\rm 1\hskip-3.87495pt1}_{\{|N_{k}|>\varepsilon b_{n}/2\}})].$
By Chebyshev’s association inequality we deduce
$\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]$ $\leq\frac{4}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|Y_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|>\varepsilon b_{n}/2\}}]+\frac{4}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|N_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|N_{k}|>\varepsilon b_{n}/2\}}].$
It is easy to check that $\frac{4}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|N_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|N_{k}|>\varepsilon b_{n}/2\}}]\to 0$ as $n\to\infty.$ Hence, by the condition (1.7), we conclude that the sequence $(X_{k})_{k\geq 1}$ also satisfies Lindeberg’s condition.
On the other hand, by Lemma 1.20 in [9], the score function of $X_{k}$ is given by
$\rho_{k}(X_{k})=\mathbb{E}\left[\frac{-N_{k}}{\alpha_{k}^{2}}\bigg|X_{k}\right],\,\,k\geq 1.$
This implies that $\sup\limits_{k\geq 1}J(X_{k})<\infty.$ Indeed, we have
$J(X_{k})={\rm Var}(X_{k})\mathbb{E}[\rho_{k}^{2}(X_{k})]=2\alpha_{k}^{2}\mathbb{E}\left(\mathbb{E}\left[\frac{-N_{k}}{\alpha_{k}^{2}}\bigg|X_{k}\right]\right)^{2}\leq 2,\,\,k\geq 1.$
Thanks to Corollary 1.1, we obtain $d_{TV}(S_{n},N)\to 0$ as $n\to\infty.$ For every $t\in\mathbb{R},$ in view of Lévy’s continuity theorem, we have
$\mathbb{E}\left[\exp\left(it\frac{Y_{1}+\cdots+Y_{n}}{a_{n}}\right)\right]$ $=e^{t^{2}/2}\mathbb{E}\left[\exp\left(it\frac{Y_{1}+N_{1}+\cdots+Y_{n}+N_{n}}{a_{n}}\right)\right]$
$=e^{t^{2}/2}\mathbb{E}\left[\exp\left(it\sqrt{2}S_{n}\right)\right]\to e^{-t^{2}/2},\,\,n\to\infty$
and hence, $\frac{Y_{1}+\cdots+Y_{n}}{a_{n}}\to N$ in distribution.
The proof is complete. $\square$